A block is kept on the floor of an elevator at reat.
The elevator start decending with an acceleration
Of 12m/sec2.Find the displacement of the block during the first 0.2 second after the start. Take g=10m/sec2

Grade:12th pass

1 Answers

Khimraj
4 years ago
Since the acceleration of elevator is greater than g (Acceleration due to gravity 9.8 m/s2 ) and block and elevator are not attached, the block will get separated and move like free fall body with acceleration g. we know s=ut+1/2at^2 where u=0 and a=g(free falll acceleration) s=1/2 * 10* 0.2^2...